### 13 Mar 2026 v1

# warped area minimal hypersurface and warped product metric

Abstract.
By studying the warped(or weighted) area-minimizing hypersurface, we prove that the metric can be locally split as a warped product metric under the spectral Ricci or spectral scalar curvature lower bound condition.
Key words and phrases:
warped product metric, area-minimizing hypersurface, spectral curvature bound
2020 Mathematics Subject Classification:
53C24
1. Introduction
A classic rigidity result concerning area-minimizing hypersurfaces, due to Cai-Galloway[CG00], asserts the following: if a three-dimensional manifold $(M,g)$ has non-negative scalar curvature ($\operatorname{Sc}_{g}\geq 0$) and contains a two-sided, locally area-minimizing torus $\Sigma$, then $M$ is flat in a neighborhood of $\Sigma$; equivalently, $M$ admits a product metric near $\Sigma$. This confirms a conjecture originally proposed by Fischer-Colbrie-Schoen[FCS80]. Building on the results of Schoen-Yau[SY79] on the classifications of area-minimizing surfaces in three-dimensional manifolds with positive scalar curvature, any such surface in $M$ must be homeomorphic to either $\mathbb{S}^{2}$ (the 2-sphere) or $\mathbb{RP}^{2}$ (the real projective plane). Subsequently, Bray-Brendle-Neves[BBN10] investigated the case of an area-minimizing 2-sphere $\Sigma$ in a compact three-dimensional manifold $M^{3}$ with scalar curvature bounded below by 2. In particular, they proved that the area of $\Sigma$ is at most $4\pi$; crucially, equality holds if and only if $M^{3}$ is covered by $\mathbb{S}^{2}\times\mathbb{R}$. This implies that, in the rigid case, $M$ locally admits a product metric near $\Sigma$. Since then, additional rigidity results establishing local product metric splittings for area-minimizing hypersurfaces have been obtained [CEM19],[Zhu20],[CLZ25].
A natural question thus arises: under what conditions do hypersurfaces admit local splittings as warped product metrics? In this paper, we address this question and establish several key results. We further elucidate the geometric significance of warped(or weighted) area-minimizing hypersurfaces. To begin with, we analyze the Ricci curvature and scalar curvature of the warped product manifold $(\mathbb{S}^{1}\times N^{n-1},g=dt^{2}+f^{2}(t)g_{N})$ where $f$ denotes a positive smooth function on $\mathbb{S}^{1}$ and $(N^{n-1},g_{N})$ is a Riemannian manifold. The Ricci curvature of the metric $g$ is
(1.1) $\operatorname{Ric}_{g}(\partial_{t},\partial_{t})=$ $-(n-1)\frac{f^{\prime\prime}(t)}{f(t)},$
(1.2) $\operatorname{Ric}_{g}(X,Y)=$ $\operatorname{Ric}_{g_{N}}(X,Y)-\left[\frac{f^{\prime\prime}(t)}{f(t)}+(n-2)\frac{(f^{\prime}(t))^{2}}{f^{2}(t)}\right]\langle X,Y\rangle_{g},$
where $X,Y\in TN$. The scalar curvature of the metric $g$ is
(1.3) $\operatorname{Sc}_{g}=$ $\frac{\operatorname{Sc}_{g_{N}}}{f^{2}}-2(n-1)\frac{f^{\prime\prime}(t)}{f(t)}-(n-1)(n-2)\frac{(f^{\prime}(t))^{2}}{f^{2}}.$
Then, by direct computation, we have
(1.4) $-(n-1)f\Delta_{g}f^{-1}+\operatorname{Ric}_{g}(\partial_{t},\partial_{t})=$ $(n-1)(n-3)f^{-2}(f^{\prime})^{2}$
(1.5) $\quad-(n-1)f\Delta_{g}f^{-1}+\frac{1}{2}\operatorname{Sc}_{g}=$ $\frac{(n-1)(n-4)}{2}f^{-2}(f^{\prime})^{2}\text{, if }\operatorname{Sc}_{g_{N}}=0.$
The warped(or weighted) volume
$E(N^{n-1})=\int_{N}f^{-(n-1)}d\operatorname{vol}_{g|_{N}}=\operatorname{Vol}g_{N}(N)$
is constant. Hence, the first and second variations of $E(N^{n-1}_{t})$ along the $\partial_{t}$ direction are all zero.
Therefore, we study the warped minimal hypersurface under the spectral Ricci curvature condition and the spectral scalar curvature condition.
Definition 1.1.
Let $(M,g)$ be a Riemannian manifold and $u$ be a positive smooth function on $M$, the spectral scalar curvature is defined as
(1.6) $\gamma u^{-1}\Delta_{g}u+\tfrac{1}{2}\operatorname{Sc}_{g}$
and the spectral Ricci curvature is defined as
(1.7) $\gamma u^{-1}\Delta_{g}u+\operatorname{Ric}_{g}$
where
(1.8) $\operatorname{Ric}_{g}:=\inf_{e\in T_{x}M,|e|_{g}=1}\operatorname{Ric}_{g}(e,e)$
is the least Ricci curvature at a given point $x\in M$.
Throughout this paper, we let $\gamma=n-1$.
Theorem 1.2.
Let $(M^{n},g)$ be a smooth compact connected Riemannian manifold satisfying
(1.9) $-\gamma u^{-1}\Delta_{g}u+\operatorname{Ric}_{g}\geq(n-1)(n-3)u^{-2}|\nabla u|^{2},$
where $u$ is a smooth positive function on $M$ and $3\leq n\leq 7$. Suppose further that $H_{n-1}(M,\mathbb{Z})\neq 0$. Then there exists a non-trivial homology class $\Sigma]\in H_{n-1}(M,\mathbb{Z})$(i.e., $\Sigma]\neq 0$) such that:
  1. (1)
    $u$ is constant on $\Sigma$;
  2. (2)
    $(M^{n},g)$ locally splits as $(M^{n},g=dt^{2}+f^{2}(t)g_{\Sigma})$ with $f(t)=u^{-1}(t)$;
  3. (3)
    the Ricci curvature of $\Sigma$ satisfies
    $\operatorname{Ric}_{g_{\Sigma}}(\cdot,\cdot)\geqslant(n-2)\left[(f^{\prime}(t))^{2}-f(t)f^{\prime\prime}(t)\right]g_{\Sigma}(\cdot,\cdot).$
Remark 1.3.
For $n=3$, then $2u^{-1}\Delta_{g}u+\operatorname{Ric}_{g}\geq 0$, i.e., the spectral Ricci curvature is nonnegative. For a Riemannian manifold $(M^{n},g)$ with two ends satisfying the nonnegative spectral Ricci curvature
$a\Delta_{g}u+\operatorname{Ric}_{g}u\geq 0$
with $a<\frac{4}{n-1}$. Catino-Mari-Mastrolia-Roncoroni[CMMR25] and Antonelli-Pozzetta-Xu[APX24] proved that $M$ splits as a product manifold. As Antonelli-Pozzetta-Xu pointed out, such a product metric splitting fails to hold when $a\geq\frac{4}{n-1}$. Nevertheless, for the case $n=3$, if there exists a compact warped area-minimizing hypersurface, then the manifold $M^{3}$ with two ends will locally split as a warped product metric according to Theorem 1.2. For nonnegative Bakry–Émery Ricci curvature, Liu [Liu13] obtains a locally split product metric by studying the weighted area-minimizing hypersurface.
For the spectral scalar curvature, we have
Theorem 1.4.
Let $(M^{n},g)$ be a smooth compact connected Riemannian manifold satisfying
(1.10) $\gamma u^{-1}\Delta_{g}u+\frac{1}{2}\operatorname{Sc}_{g}\geq\frac{\gamma(\gamma-3)}{2}u^{-2}|\nabla u|^{2},$
where $u$ is a smooth positive function on $M$ and $5\leq n\leq 7$. If $H_{n-1}(M,\mathbb{Z})\neq 0$ and $0\neq[\Sigma]\in H_{n-1}(M,\mathbb{Z})$ such that there is no positive scalar curvature on $\Sigma$, then, $u$ is constant on $\Sigma$ and $(M^{n},g)$ locally splits as $(M^{n},g=dt^{2}+u^{-2}g_{\Sigma})$. Moreover, $(\Sigma,g_{\Sigma})$ is Ricci flat.
[…]
Remark 1.5.
The operator $\gamma\Delta_{g}+\frac{1}{2}\operatorname{Sc}_{g}$ is greater than the conformal operator $\frac{2(n-1)}{n-2}\Delta_{g}+\frac{1}{2}\operatorname{Sc}_{g}$ for $5\leq n\leq 7$. That is, if the $\gamma\Delta_{g}+\frac{1}{2}\operatorname{Sc}_{g}$ is positive, we can not use the conformal change to obtain a positive scalar curvature metric.
The proofs of Theorem 1.2 and Theorem 1.4 rely on the study of warped(or weighted) area-minimizing hypersurfaces. Specifically, we analyze the first and the second variations of such hypersurfaces, then employ a foliation argument analogous to that used in the theory of area-minimizing hypersurfaces [BBN10, CS25] to establish the theorems.
The organization of this paper is as follows: In section 2, we compute the first and second variation of warped(or weighted) area-minimizing hypersurfaces. In section 3, we present the proof of Theorem 1.2. In section 4, we prove Theorem 1.4.
2. Basic fact about warped or weighted minimal hypersurface
Let $(M^{n},g)$ be a compact smooth Riemannian manifold, $\Sigma$ a hypersurface embedded in $M^{n}$, and $u\in C^{\infty}(M)$ a positive function. We define the warped(or weighted) volume of $\Sigma$ by the integral
(2.1) $E(\Sigma)=\int_{\Sigma}u^{\gamma}\mathrm{d}\mathcal{H}^{n-1},$
where $\gamma=n-1$ and $\mathcal{H}^{n-1}$ denotes the (n1)(n-1) dimension Hausdorff measure induced by the metric $g$. Let $H$ and $A$ stand for the mean curvature and the second fundamental of $\Sigma$, respectively.
Lemma 2.1.
Let $\Sigma_{t}$ be a variation of $\Sigma$ with the variation vector field given by $Y$ and $\nu$ be the unit normal vector of $\Sigma$, then
(2.2) $\tfrac{\mathrm{d}}{\mathrm{d}t}E(\Sigma_{t})|_{t=0}=\int_{\Sigma}(H+\gamma u^{-1}u_{\nu})\langle Y,\nu\rangle u^{\gamma}\mathrm{d}\mathcal{H}^{n-1}.$
We say $\Sigma$ a warped(or weighted) area-minimizing hypersurface if $H+\gamma u^{-1}u_{\nu}=0$ along $\Sigma$.
Lemma 2.2.
Consider a warped area-minimizing hypersurface $\Sigma$. Assume that $\Sigma_{t}$ is a smooth $1$-parameter family of smooth hypersurfaces along the variation vector field $Y$ and normal speed $\phi=\langle Y,\nu\rangle$ at $t=0$. Then
(2.3) $\tfrac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}E(\Sigma_{t})|_{t=0}=$ $\int_{\Sigma}[-\Delta_{\partial\Omega}\phi-|A|^{2}\phi-\operatorname{Ric}(\nu,\nu)\phi-\gamma u^{-2}u_{\nu}^{2}\phi$
(2.4) $\qquad+\gamma u^{-1}\phi(\Delta u-\Delta_{\partial\Omega}u-Hu_{\nu})-\gamma u^{-1}\langle\nabla_{\partial\Omega}u,\nabla_{\partial\Omega}\phi\rangle]u^{\gamma}\phi.$
A warped area-minimizing hypersurface $\Sigma$ is called stable if
$\tfrac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}E(\Sigma_{t})|_{t=0}\geq 0.$
Lemma 2.3.
The second variation (2.4) of the functional $E(\Sigma_{t})$ in above can be rewritten as
$\tfrac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}E(\Sigma_{t})|_{t=0}$ $=\int_{\Sigma}|\nabla_{\Sigma}\psi|^{2}+\int_{\Sigma}(\gamma\psi\langle\nabla_{\Sigma}w,\nabla_{\Sigma}\psi\rangle+(\tfrac{\gamma^{2}}{4}-\gamma)\psi^{2}|\nabla_{\Sigma}w|^{2})$
$\quad+\int_{\Sigma}(\gamma u^{-1}\Delta_{g}u-(|A|^{2}+\operatorname{Ric}(\nu,\nu)))\psi^{2}$
$\quad-\int_{\Sigma}(\gamma Hw_{\nu}+\gamma w_{\nu}^{2})\psi^{2}.$
where $\psi=\phi u^{\gamma/2}$ and $w=\log u$.
Proof.
See [CS25, Lemma 2.4]. ∎
3. Spectral Ricci case
In this section, we prove Theorem 1.2. We first analyze the first and second variations to obtain some key inequalities.
Proof of Theorem 1.2.
Since $H_{n-1}(M,\mathbb{Z})\neq 0$, by geometric theorem(Section 5.1.6[Fed69]), there exists a stable warped area-minimizing hypersurface in $\Sigma$. By Lemma 2.3, we have
$0\leq$ $\int_{\Sigma}|\nabla_{\Sigma}\psi|^{2}+\int_{\Sigma}(\gamma\psi\langle\nabla_{\Sigma}w,\nabla_{\Sigma}\psi\rangle+(\tfrac{\gamma^{2}}{4}-\gamma)\psi^{2}|\nabla_{\Sigma}w|^{2})$
$\quad+\int_{\Sigma}(-\gamma(n-3)u^{-2}|\nabla u|^{2})\psi^{2}$
$\quad-\int_{\Sigma}\left(\frac{1}{n-1}H^{2}+\gamma Hw_{\nu}+\gamma w_{\nu}^{2}\right)\psi^{2}$
$\leq$ $\int_{\Sigma}|\nabla_{\Sigma}\psi|^{2}+\int_{\Sigma}(\gamma\psi\langle\nabla_{\Sigma}w,\nabla_{\Sigma}\psi\rangle+\left(\tfrac{\gamma}{4}-n+2\right)\gamma\psi^{2}|\nabla_{\Sigma}w|^{2})$
$\quad-\int_{\Sigma}\left(\frac{1}{n-1}H^{2}+\gamma Hw_{\nu}+\gamma(n-2)w_{\nu}^{2}\right)\psi^{2},$
Here we have used $u^{-2}|\nabla u|^{2}=|\nabla w|^{2}=|\nabla_{\Sigma}w|^{2}+|w_{\nu}|^{2}$.
From the following equality,
$\gamma\psi\langle\nabla_{\Sigma}w,\nabla_{\Sigma}\psi\rangle+\left(\tfrac{\gamma}{4}-n+2\right)\gamma\psi^{2}|\nabla_{\Sigma}w|^{2}$
$=$ $\gamma\left(\tfrac{\gamma}{4}-n+2\right)\left|\psi\nabla_{\Sigma}w+\frac{1}{2\left(\tfrac{\gamma}{4}-n+2\right)}\nabla_{\Sigma}\psi\right|^{2}$
$-\frac{\gamma}{4\left(\tfrac{\gamma}{4}-n+2\right)}|\nabla_{\Sigma}\psi|^{2},$
we have
$0\leq$ $\int_{\Sigma}|\nabla_{\Sigma}\psi|^{2}+\int_{\Sigma}(\gamma\psi\langle\nabla_{\Sigma}w,\nabla_{\Sigma}\psi\rangle+\left(\tfrac{\gamma}{4}-n+2\right)\gamma\psi^{2}|\nabla_{\Sigma}w|^{2})$
$\quad-\int_{\Sigma}(\frac{1}{n-1}H^{2}+\gamma Hw_{\nu}+\gamma(n-2)w_{\nu}^{2})\psi^{2}$
$=$ $\int_{\Sigma}\frac{4n-8}{4n-8-\gamma}|\nabla_{\Sigma}\psi|^{2}+\int_{\Sigma}\gamma\left(\tfrac{\gamma}{4}-n+2\right)\left|\psi\nabla_{\Sigma}w+\frac{1}{2\left(\tfrac{\gamma}{4}-n+2\right)}\nabla_{\Sigma}\psi\right|^{2}$
$\quad-\int_{\Sigma}\left(\frac{1}{n-1}H^{2}+\gamma Hw_{\nu}+\gamma(n-2)w_{\nu}^{2}\right)\psi^{2}.$
Hence, by $H+\gamma w_{\nu}=0$, we have
$\int_{\Sigma}\frac{4n-8}{3n-7}|\nabla_{\Sigma}\psi|^{2}\geq$ $\int_{\Sigma}\frac{(n-1)\left(3n-7\right)}{4}\left|\psi\nabla_{\Sigma}w+\frac{2}{\left(7-3n\right)}\nabla_{\Sigma}\psi\right|^{2}$
$\quad+\int_{\Sigma}\left(\frac{1}{n-1}H^{2}+(n-1)Hw_{\nu}+(n-1)(n-2)w_{\nu}^{2}\right)\psi^{2}$
$=$ $\int_{\Sigma}\frac{(n-1)\left(3n-7\right)}{4}\left|\psi\nabla_{\Sigma}w+\frac{2}{\left(7-3n\right)}\nabla_{\Sigma}\psi\right|^{2}.$
Taking $\psi=1$, we can obtain $\nabla_{\Sigma}w=0$, $A=\frac{H}{n-1}g|_{\Sigma}$, $\operatorname{Ric}_{g}(\nu,\nu)=\operatorname{Ric}_{g}$ and $\gamma u^{-1}\Delta_{g}u+\operatorname{Ric}_{g}=(n-1)(n-3)w_{\nu}^{2}$ on $\Sigma$. The remaining proof is using the foliation argument. We do this in Lemma 3.1, Lemma 3.2
The following argument is standard(see [BBN10],[Zhu21],[CS25]).
Lemma 3.1.
Let $\Sigma$ be a stable warped area-minimizing hypersurface. We can construct a local foliation $\{\Sigma_{t}\}_{-\epsilon<t<\epsilon}$ such that $\Sigma_{t}$ is of constant $H_{t}+\gamma w_{\nu_{t}}$, $\Sigma_{0}=\Sigma$,
  1. (1)
    each $\Sigma_{t}$ is a graph over $\Sigma$ with graph function $\rho_{t}$ along outward unit normal vector field $\nu$ such that
    (3.1) $\left.\frac{\partial\rho_{t}}{\partial t}\right|_{t=0}=1\quad\text{ and }\quad\frac{1}{\operatorname{vol}(\Sigma)}\int_{\Sigma}\rho_{t}dv=t;$
  2. (2)
    and $H_{t}+\gamma\omega_{\nu_{t}}$ is constant on $\Sigma_{t}$.
Here, $dv$ is the volume element of $\Sigma$ and $H_{t}$ is the mean curvature of $\Sigma_{t}$.
Proof.
Let
$\hat{C}^{\alpha}(\Sigma)=\left\{\phi\in C^{\alpha}(\Sigma):\text{ }\int_{\Sigma}\phi dv=0\right\}$
for some $\alpha\in(0,1)$. Consider the map
$\tilde{\Phi}:C^{2,\alpha}(\Sigma)$ $\to$ $\hat{C}^{\alpha}(\Sigma)\times\mathbb{R},$
$\rho$ $\mapsto$ $\left(H_{\rho}+\gamma\omega_{\nu}-\frac{1}{\operatorname{vol}(\Sigma)}{\int_{\Sigma}\left(H_{\rho}+\gamma\omega_{\nu}\right)dv},\frac{1}{\operatorname{vol}(\Sigma)}{\int_{\Sigma}\rho dv}\right),$
where $H_{\rho}$ is the mean curvature of the graph over $\Sigma$ with graph function $\rho$.
We calculate the first variation of $H+\gamma u^{-1}u_{\nu}$ at $t=0$:
$\delta_{\phi\nu}(H+\gamma u^{-1}u_{\nu})$
$=$ $-\Delta_{\Sigma}\phi-(\operatorname{Ric}(\nu,\nu)+|A|^{2})\phi-\gamma u^{-2}u_{\nu}^{2}\phi-\gamma u^{-1}\langle\nabla_{\Sigma}u,\nabla_{\Sigma}\phi\rangle$
$\quad+\gamma u^{-1}(\Delta u-\Delta_{\Sigma}u-Hu_{\nu})\phi$
$=$ $-\Delta_{\Sigma}\phi-\operatorname{Ric}(\nu,\nu)\phi-|A|^{2}\phi-\gamma w_{\nu}^{2}\phi$
$\quad+\gamma u^{-1}\phi\Delta u-\gamma u^{-1}\phi\Delta_{\Sigma}u-\gamma w_{\nu}(-\gamma w_{\nu})\phi$
$=$ $-\Delta_{\Sigma}\phi-\gamma(n-3)w_{\nu}^{2}\phi+\frac{n-2}{n-1}\gamma^{2}(w_{\nu})^{2}\phi-\gamma w_{\nu}^{2}\phi$
$=$ $-\Delta_{\Sigma}\phi.$
Then, the linearization of $\tilde{\Phi}$ at $\rho=0$ given by
(3.2) $D\tilde{\Phi}|_{\rho=0}:C^{2,\alpha}(\Sigma)\to\hat{C}^{\alpha}(\Sigma)\times\mathbb{R},\quad\psi\mapsto\left(-\Delta_{\Sigma}\psi,\frac{1}{\operatorname{vol}(\Sigma)}\int_{\Sigma}\psi dv\right),$
is invertible. By the inverse function theorem, we can find a family of functions $\rho_{t}:\Sigma\to\mathbb{R}$ with $t\in(-\epsilon,\epsilon)$ with the following properties:
  1. (1)
    the function $\rho_{t}$ satisfies $\rho_{0}\equiv 0$,
    (3.3) $\left.\frac{\partial\rho_{t}}{\partial t}\right|_{t=0}\equiv 1,\quad\text{and}\quad\frac{1}{\operatorname{vol}(\Sigma)}\int_{\Sigma}\rho_{t}dv=t$
  2. (2)
    the graphs $\Sigma_{t}$ over $\Sigma$ with the graph function $\rho_{t}$ is of constant $H+\gamma u^{-1}u_{\nu}$.
From (3.3), with the value of $\epsilon$ decreased a little bit, the speed $\partial_{t}\rho_{t}$ will be positive for $t\in(-\epsilon,\epsilon)$, from which it follows that the graphs $\{\Sigma_{t}\}_{t\in(-\epsilon,\epsilon)}$ form a foliation around $\Sigma$. ∎
Lemma 3.2.
$\Sigma_{t}$ is also a stable warped area-minimizing hypersurface in $M$.
Proof.
The first step is to prove that $H_{t}+\gamma\omega_{\nu}=0$ on $\Sigma_{t}$ for $t\in(-\epsilon,\epsilon)$. Then the stable warped area-minimizing property of $\Sigma_{t}$ follows. We do that in Proposition 3.3 and 3.4. ∎
Proposition 3.3.
There exists a continuous function $\Psi(t)$ such that
(3.4) $\tfrac{\mathrm{d}}{\mathrm{d}t}\left(\exp\left(\int_{0}^{t}\Psi(\tau)\mathrm{d}\tau\right)\tilde{H}\right)\leq 0$
where
(3.5) $\tilde{H}(t)=H_{t}+\gamma w_{\nu_{t}}.$
Proof.
Let $\Phi:\Sigma\times(-\varepsilon,\varepsilon)\to M$ parametrize the local foliation, $Y=\tfrac{\partial\Phi}{\partial t}$, and $\phi_{t}=\langle Y,\nu_{t}\rangle$. Since we have shown that $\phi_{0}$ is a constant. We can fix $\varepsilon$ sufficiently small so that $\phi_{t}>0$ for all $t\in(-\varepsilon,\varepsilon)$. Recall that the first variation gives
$-\tilde{H}^{\prime}(t)$
$=$ $-\tfrac{\mathrm{d}}{\mathrm{d}t}(H_{t}+\gamma w_{\nu_{t}})$
$=$ $\Delta_{\Sigma_{t}}\phi_{t}+(\operatorname{Ric}(\nu_{t},\nu_{t})+|A_{t}|^{2})\phi_{t}+\gamma u^{-2}u_{\nu_{t}}^{2}\phi_{t}$
$\quad-\gamma u^{-1}(\Delta_{g}u-\Delta_{\Sigma_{t}}u-H_{t}u_{\nu_{t}})\phi_{t}+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle.$
Using the definition of $\operatorname{Ric}$, (4.4), and the spectral bound, and with suitable grouping of terms, we see
$-\tilde{H}^{\prime}(t)\phi_{t}^{-1}$
$\geq$ $(\phi_{t}^{-1}\Delta_{\Sigma_{t}}\phi_{t}+\gamma u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\phi_{t}^{-1})$
$\quad+\frac{1}{n-1}H_{t}^{2}+\gamma w_{\nu_{t}}^{2}+\gamma H_{t}w_{\nu_{t}}+\gamma(n-3)u^{-2}|\nabla u|^{2}.$
Inserting (4.4) in the above and using
$(\phi_{t}^{-1}\Delta_{\Sigma_{t}}\phi_{t}+\gamma u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\phi_{t}^{-1})$
$=$ $\operatorname{div}_{\Sigma_{t}}\left(\frac{\nabla_{\Sigma_{t}}\phi_{t}}{\phi_{t}}+\gamma\frac{\nabla_{\Sigma_{t}}u}{u}\right)+(1-\tfrac{\gamma}{4})\left|\tfrac{\nabla_{\Sigma_{t}}\phi_{t}}{\phi_{t}}\right|^{2}+\gamma\left|\frac{\nabla_{\Sigma_{t}}u}{u}-\frac{\nabla_{\Sigma_{t}}\phi_{t}}{2\phi_{t}}\right|^{2}$
$\geq$ $\operatorname{div}_{\Sigma_{t}}\left(\frac{\nabla_{\Sigma_{t}}\phi_{t}}{\phi_{t}}+\gamma\frac{\nabla_{\Sigma_{t}}u}{u}\right)$
yields
$-\tilde{H}^{\prime}(t)\phi_{t}^{-1}$
$\geq$ $\operatorname{div}_{\Sigma_{t}}\left(\frac{\nabla_{\Sigma_{t}}\phi_{t}}{\phi_{t}}+\gamma\frac{\nabla_{\Sigma_{t}}u}{u}\right)+\gamma(n-3)u^{-2}|\nabla_{\Sigma_{t}}u|^{2}+\gamma(n-3)u^{-2}u_{\nu_{t}}^{2}$
$\quad+\frac{1}{n-1}\tilde{H}^{2}-\frac{2}{n-1}\tilde{H}\gamma w_{\nu_{t}}+\frac{1}{n-1}\gamma^{2}w_{\nu_{t}}^{2}+\gamma w_{\nu_{t}}^{2}+\gamma\tilde{H}w_{\nu_{t}}-\gamma^{2}w_{\nu_{t}}^{2}$
$\geq$ $\operatorname{div}_{\Sigma_{t}}\left(\frac{\nabla_{\Sigma_{t}}\phi_{t}}{\phi_{t}}+\gamma\frac{\nabla_{\Sigma_{t}}u}{u}\right)$
$\quad+\frac{1}{n-1}\tilde{H}^{2}+\frac{n-3}{n-1}\tilde{H}\gamma w_{\nu_{t}}+\left(\frac{2-n}{n-1}\gamma+n-2\right)\gamma w_{\nu_{t}}^{2}$
By the trivial bound $\tilde{H}^{2}\geq 0$, we obtain that
$-\tilde{H}^{\prime}(t)\phi_{t}^{-1}$
$\geq$ $\operatorname{div}_{\Sigma_{t}}\left(\frac{\nabla_{\Sigma_{t}}\phi_{t}}{\phi_{t}}+\gamma\frac{\nabla_{\Sigma_{t}}u}{u}\right)+\frac{n-3}{n-1}\tilde{H}\gamma w_{\nu_{t}}.$
We integrate the above on $\Sigma_{t}$, and we find by the divergence theorem that,
$\tilde{H}^{\prime}(t)\int_{\Sigma_{t}}\tfrac{1}{\phi_{t}}\geq\tilde{H}(t)\int_{\Sigma_{t}}(n-3)w_{\nu_{t}}.$
We set $\Psi(t)=(\int_{\Sigma_{t}}\tfrac{1}{\phi_{t}})^{-1}\int_{\Sigma_{t}}(n-3)w_{\nu_{t}}$, then
$\tilde{H}^{\prime}+\Psi(t)\tilde{H}\leq 0.$
By solving this inequality, we finish the proof of the Proposition. ∎
Proposition 3.4.
Every $\Sigma_{t}$ is a stable warped area-minimizing in $M$.
Proof.
Let $\partial_{t}$ be the variation vector field of the foliation $\{\Sigma_{t}\}_{t\in(-\varepsilon,\varepsilon)}$, and $\phi_{t}=\langle\partial_{t},\nu_{t}\rangle$. Recall that the first variation $E(\Sigma_{t})$
$\tfrac{\mathrm{d}}{\mathrm{d}t}E(\Sigma_{t})=\int_{\Sigma_{t}}\tilde{H}(t)u^{\gamma}\phi_{t}.$
It follows from $\tilde{H}(0)=0$ and (3.4) that $\tilde{H}(t)\leq 0$ for $t\geq 0$ and $\tilde{H}(t)\geq 0$ for $t\leq 0$. So $E(\Sigma_{t})\leq E(\Sigma_{0})$ for all $t\in(-\varepsilon,\varepsilon)$ and hence
$E(\Sigma_{t})=E(\Sigma_{0})$
for all $t\in(-\varepsilon,\varepsilon)$. Therefore, all the foliation analysis on $\Sigma_{0}$ can be applied to $\Sigma_{t}$. Since $M$ is connected, we can conclude that $M$ is foliated by the stable warped area-minimizing hypersurface $\Sigma_{t}$. ∎
In the following, we deduce that the metric $g$ can be locally written as a warped product metric. From lemma 3.1 and lemma 3.2, we can obtain that, locally, $M^{n}$ can be written as $(-\epsilon,\epsilon)\times\Sigma$ such that each $\Sigma_{t}$ is a warped area-minimizing hypersurface and $\Sigma_{t}$ is umbilical. Then the metric $g$ can be written as $g=dt^{2}+f^{2}(t)g_{\Sigma}$ for some positive function $f$ on $(-\epsilon,\epsilon)$. Since $H_{t}=-(n-1)u^{-1}u^{\prime}(t)=(n-1)f^{-1}f^{\prime}(t)$, we obtain $f(t)=u^{-1}(t)$. From equations (1.1) and (1.2), we have
$\operatorname{Ric}_{g_{\Sigma}}(\cdot,\cdot)\geq(n-2)[(f^{\prime}(t))^{2}-f(t)f^{\prime\prime}(t)]g_{\Sigma}(\cdot,\cdot)$
By an open and closed argument, passing to a suitable cover $\tilde{M}$, $(\tilde{M}^{n},g)$ is isometric to $(\mathbb{S}^{1}\times\Sigma^{n-1},ds^{2}+f^{2}(t)g|_{\Sigma})$.
4. Spectral scalar curvature case
In this section, we prove Theorem 1.4. The proof is analogous to the case of the spectral Ricci case. However, the details are different.
Proof.
Since $H_{n-1}(M,\mathbb{Z})\neq 0$, by the geometric theorem, there exists a stable warped area-minimizing hypersurface in $\Sigma$. Using Schoen-Yau’s trick, rewrite the Gauss equations as
$|A|^{2}+\operatorname{Ric}(\nu,\nu)=\frac{1}{2}(|A|^{2}+H^{2}+\operatorname{Sc}_{g}-\operatorname{Sc}_{\Sigma}).$
By Lemma 2.3, we have
$0\leq$ $\int_{\Sigma}|\nabla_{\Sigma}\psi|^{2}+\int_{\Sigma}(\gamma\psi\langle\nabla_{\Sigma}w,\nabla_{\Sigma}\psi\rangle+(\tfrac{\gamma^{2}}{4}-\gamma)\psi^{2}|\nabla_{\Sigma}w|^{2})$
$\quad+\int_{\Sigma}(\gamma u^{-1}\Delta_{g}u-(|A|^{2}+\operatorname{Ric}(\nu,\nu)))\psi^{2}$
$\quad-\int_{\Sigma}(\gamma Hw_{\nu}+\gamma w_{\nu}^{2})\psi^{2}$
$=$ $\int_{\Sigma}|\nabla_{\Sigma}\psi|^{2}+\int_{\Sigma}(\gamma\psi\langle\nabla_{\Sigma}w,\nabla_{\Sigma}\psi\rangle+(\tfrac{\gamma^{2}}{4}-\gamma)\psi^{2}|\nabla_{\Sigma}w|^{2})$
$\quad+\int_{\Sigma}(\gamma u^{-1}\Delta_{g}u-\frac{1}{2}(|A|^{2}+H^{2}+\operatorname{Sc}_{g}-\operatorname{Sc}_{\Sigma}))\psi^{2}$
$\quad-\int_{\Sigma}(\gamma Hw_{\nu}+\gamma w_{\nu}^{2})\psi^{2}.$
Recall $\gamma=n-1$. By $H=-\gamma w_{\nu}$, we obtain that
$0\leq$ $\int_{\Sigma}(|\nabla_{\Sigma}\psi|^{2}+\frac{1}{2}\operatorname{Sc}_{\Sigma}\psi^{2})+\int_{\Sigma}(\gamma\psi\langle\nabla_{\Sigma}w,\nabla_{\Sigma}\psi\rangle+(\tfrac{\gamma^{2}}{4}-\gamma)\psi^{2}|\nabla_{\Sigma}w|^{2})$
$\quad-\int_{\Sigma}\left(\frac{\gamma(\gamma-3)}{2}u^{-2}|\nabla u|^{2}+\frac{n}{2(n-1)}H^{2}\right)\psi^{2}$
$\quad-\int_{\Sigma}(\gamma Hw_{\nu}+\gamma w_{\nu}^{2})\psi^{2}$
$=$ $\int_{\Sigma}\left(|\nabla_{\Sigma}\psi|^{2}+\frac{1}{2}\operatorname{Sc}_{\Sigma}\psi^{2}\right)+\int_{\Sigma}(\gamma\psi\langle\nabla_{\Sigma}w,\nabla_{\Sigma}\psi\rangle+\tfrac{\gamma(2-\gamma)}{4}\psi^{2}|\nabla_{\Sigma}w|^{2})$
$\quad-\int_{\Sigma}\left(\frac{\gamma(\gamma-3)}{2}w_{\nu}^{2}+\frac{n}{2(n-1)}H^{2}+\gamma Hw_{\nu}+\gamma w_{\nu}^{2}\right)\psi^{2}$
$=$ $\int_{\Sigma}\left(|\nabla_{\Sigma}\psi|^{2}+\frac{1}{2}\operatorname{Sc}_{\Sigma}\psi^{2}\right)+\int_{\Sigma}(\gamma\psi\langle\nabla_{\Sigma}w,\nabla_{\Sigma}\psi\rangle+\tfrac{\gamma(2-\gamma)}{4}\psi^{2}|\nabla_{\Sigma}w|^{2}),$
where we have used that $u^{-2}|\nabla u|^{2}=|\nabla w|^{2}=|\nabla_{\Sigma}w|^{2}+|w_{\nu}|^{2}$. Since
$\frac{\gamma(2-\gamma)}{4}\psi^{2}|\nabla_{\Sigma}w|^{2}+\gamma\psi\langle\nabla_{\Sigma}w,\nabla_{\Sigma}\psi\rangle$
$=$ $\frac{\gamma(2-\gamma)}{4}\left|\psi\nabla_{\Sigma}w+\frac{2}{2-\gamma}\nabla_{\Sigma}\psi\right|^{2}-\frac{\gamma}{2-\gamma}|\nabla_{\Sigma}\psi|^{2}.$
Then, we obtain that
$\int_{\Sigma}\left[\left(2+\frac{2}{\gamma-2}\right)|\nabla_{\Sigma}\psi|^{2}+\frac{1}{2}\operatorname{Sc}_{\Sigma}\psi^{2}\right]$
$\geq$ $\int_{\Sigma}\frac{\gamma(\gamma-2)}{4}\left|\psi\nabla_{\Sigma}w+\frac{2}{2-\gamma}\nabla_{\Sigma}\psi\right|^{2}.$
Thus, if $\nabla_{\Sigma}w\neq 0$, then the operator
$\left(2+\frac{2}{n-3}\right)\Delta_{\Sigma}+\frac{1}{2}\operatorname{Sc}_{\Sigma}=-\left(\frac{2(n-2)}{n-3}\right)\Delta_{\Sigma}+\frac{1}{2}\operatorname{Sc}_{\Sigma}$
is positive. Since $\Sigma$ admits no positive scalar curvature, by making a conformal change for the metric $g_{\Sigma}$, we will obtain a contradiction. Therefore, we have $\nabla_{\Sigma}w=0$, $\gamma u^{-1}\Delta_{g}+\frac{1}{2}\operatorname{Sc}_{g}=\frac{\gamma(\gamma-3)}{2}|\nabla w|^{2}=\frac{\gamma(\gamma-3)}{2}w_{\nu}^{2}$, $\operatorname{Sc}_{\Sigma}=0$ and $A_{\Sigma}=\frac{H}{n-1}g_{\Sigma}$. The remaining proof is using the foliation argument as in the spectral Ricci curvature case. We do this in Lemma 4.1, Lemma 4.2. ∎
Lemma 4.1.
Let $\Sigma$ be a warped area-minimizing hypersurface of $E(\Sigma)$. We can construct a local foliation $\{\Sigma_{t}\}_{-\epsilon<t<\epsilon}$ such that $\Sigma_{t}$ is of constant $H_{t}+\gamma w_{\nu_{t}}$, $\Sigma_{0}=\Sigma$,
  1. (1)
    each $\Sigma_{t}$ is a graph over $\Sigma$ with graph function $\rho_{t}$ along outward unit normal vector field $\nu$ such that
    (4.1) $\left.\frac{\partial\rho_{t}}{\partial t}\right|_{t=0}=1\quad\text{ and }\quad\frac{1}{\operatorname{vol}(\Sigma)}\int_{\Sigma}\rho_{t}dv=t;$
  2. (2)
    and $H_{t}+\gamma\omega_{\nu_{t}}$ is constant on $\Sigma_{t}$.
Here, $dv$ is the volume element of $\Sigma$ and $H_{t}$ is the mean curvature of $\Sigma_{t}$.
Proof.
Let
$\hat{C}^{\alpha}(\Sigma)=\left\{\phi\in C^{\alpha}(\Sigma):\text{ }\int_{\Sigma}\phi dv=0\right\}$
for some $\alpha\in(0,1)$. Consider the map
$\tilde{\Phi}:C^{2,\alpha}(\Sigma)$ $\to$ $\hat{C}^{\alpha}(\Sigma)\times\mathbb{R},$
$\rho$ $\mapsto$ $\left(H_{\rho}+\gamma\omega_{\nu}-\frac{1}{\operatorname{vol}(\Sigma)}{\int_{\Sigma}\left(H_{\rho}+\gamma\omega_{\nu}\right)dv},\frac{1}{\operatorname{vol}(\Sigma)}{\int_{\Sigma}\rho dv}\right),$
where $H_{\rho}$ is the mean curvature of the graph over $\Sigma$ with graph function $\rho$.
We calculate the first variation of $H+\gamma u^{-1}u_{\nu}$ at $\rho=0$:
$\delta_{\phi\nu}(H+\gamma u^{-1}u_{\nu})$
$=$ $-\Delta_{\Sigma}\phi-(\operatorname{Ric}(\nu,\nu)+|A|^{2})\phi-\gamma u^{-2}u_{\nu}^{2}\phi-\gamma u^{-1}\langle\nabla_{\Sigma}u,\nabla_{\Sigma}\phi\rangle$
$\quad+\gamma u^{-1}(\Delta u-\Delta_{\Sigma}u-Hu_{\nu})\phi$
$=$ $-\Delta_{\Sigma}\phi-\frac{1}{2}\left(\operatorname{Sc}_{g}-\operatorname{Sc}_{\Sigma}+H^{2}+|A|^{2}\right)\phi-\gamma w_{\nu}^{2}\phi$
$\quad+\gamma u^{-1}\phi\Delta u-\gamma u^{-1}\phi\Delta_{\Sigma}u-\gamma w_{\nu}(-\gamma w_{\nu})\phi$
$=$ $-\Delta_{\Sigma}\phi-\frac{\gamma(\gamma-3)}{2}w_{\nu}^{2}\phi-\frac{n}{2(n-1)}\gamma^{2}(w_{\nu})^{2}\phi-\gamma w_{\nu}^{2}\phi+\gamma^{2}w_{v}^{2}\phi$
$=$ $-\Delta_{\Sigma}\phi.$
Then, the linearization of $\tilde{\Phi}$ at $\rho=0$ given by
$D\tilde{\Phi}|_{\rho=0}:C^{2,\alpha}(\Sigma)\to\hat{C}^{\alpha}(\Sigma)\times\mathbb{R},\quad\psi\mapsto\left(-\Delta_{\Sigma}\psi,\frac{1}{\operatorname{vol}(\Sigma)}\int_{\Sigma}\psi dv\right),$
is invertible. By the inverse function theorem, we can find a family of functions $\rho_{t}:\Sigma\to\mathbb{R}$ with $t\in(-\epsilon,\epsilon)$ with the following properties:
  1. (1)
    the function $\rho_{t}$ satisfies $\rho_{0}\equiv 0$,
    (4.2) $\left.\frac{\partial\rho_{t}}{\partial t}\right|_{t=0}\equiv 1,\quad\text{and}\quad\frac{1}{\operatorname{vol}(\Sigma)}\int_{\Sigma}\rho_{t}dv=t$
  2. (2)
    the graphs $\Sigma_{t}$ over $\Sigma$ with the graph function $\rho_{t}$ is of constant $H+\gamma u^{-1}u_{\nu}$.
From (4.2), with the value of $\epsilon$ decreased a little bit, the speed $\partial_{t}\rho_{t}$ will be positive for $t\in(-\epsilon,\epsilon)$, from which it follows that the graphs $\{\Sigma_{t}\}_{t\in(-\epsilon,\epsilon)}$ form a foliation around $\Sigma$. ∎
Lemma 4.2.
$\Sigma_{t}$ is also a stable warped area-minimizing hypersurface.
Proof.
The first step is to prove that $H_{t}+\gamma\omega_{\nu_{t}}=0$ on $\Sigma_{t}$ for $t\in(-\epsilon,\epsilon)$. Then the stable warped area-minimizing property of $\Sigma_{t}$ follows. We do that in Proposition 4.3 and $4.$ ∎
Proposition 4.3.
There exists a continuous function $\Psi(t)$ such that
(4.3) $\tfrac{\mathrm{d}}{\mathrm{d}t}\left(\exp\left(\int_{0}^{t}\Psi(\tau)\mathrm{d}\tau\right)\tilde{H}\right)\leq 0$
where
(4.4) $\tilde{H}(t)=H_{t}+\gamma w_{\nu_{t}}.$
Proof.
Let $\Phi:\Sigma\times(-\varepsilon,\varepsilon)\to M$ parametrize the local foliation, $Y=\tfrac{\partial\Phi}{\partial t}$, and $\phi_{t}=\langle Y,\nu_{t}\rangle$. Since we have shown that $\phi_{0}$ is a constant. We can fix $\varepsilon$ sufficiently small so that $\phi_{t}>0$ for all $t\in(-\varepsilon,\varepsilon)$. Recall that the first variation gives
$-\tilde{H}^{\prime}(t)=$ $-\tfrac{\mathrm{d}}{\mathrm{d}t}(H_{t}+\gamma w_{\nu_{t}})$
$=$ $\Delta_{\Sigma_{t}}\phi_{t}+(\operatorname{Ric}(\nu_{t},\nu_{t})+|A_{t}|^{2})\phi_{t}+\gamma u^{-2}u_{\nu_{t}}^{2}\phi_{t}$
$\quad-\gamma u^{-1}(\Delta_{g}u-\Delta_{\Sigma_{t}}u-H_{t}u_{\nu_{t}})\phi_{t}+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle$
$=$ $\Delta_{\Sigma_{t}}\phi_{t}+\frac{1}{2}\left(\operatorname{Sc}_{g}-\operatorname{Sc}_{\Sigma_{t}}+H_{t}^{2}+|A_{t}|^{2}\right)\phi_{t}+\gamma u^{-2}u_{\nu_{t}}^{2}\phi_{t}$
$\quad-\gamma u^{-1}(\Delta_{g}u-\Delta_{\Sigma_{t}}u-H_{t}u_{\nu_{t}})\phi_{t}+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle$
Using the spectral scalar curvature bound, and with suitable grouping of terms, we see
$-\tilde{H}^{\prime}(t)\phi_{t}^{-1}$
$\geq$ $(\phi_{t}^{-1}\Delta_{\Sigma_{t}}\phi_{t}+\gamma u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\phi_{t}^{-1})-\tfrac{1}{2}\operatorname{Sc}_{\Sigma}$
$\quad+\frac{n}{2(n-1)}H_{t}^{2}+\gamma w_{\nu_{t}}^{2}+\gamma H_{t}w_{\nu_{t}}+\frac{\gamma(\gamma-3)}{2}u^{-2}|\nabla u|^{2}$
$\geq$ $(\phi_{t}^{-1}\Delta_{\Sigma_{t}}\phi_{t}+\gamma u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\phi_{t}^{-1})+\frac{\gamma(\gamma-3)}{2}u^{-2}|\nabla_{\Sigma}u|^{2}$
$\quad-\tfrac{1}{2}\operatorname{Sc}_{\Sigma}+\frac{n}{2(n-1)}\tilde{H}^{2}(t)+\tilde{H}(t)\left(\gamma w_{\nu_{t}}-\frac{n\gamma}{n-1}w_{\nu_{t}}\right)$
$+\gamma w_{\nu_{t}}^{2}+\frac{n\gamma^{2}}{2(n-1)}w_{\nu_{t}}^{2}+\frac{\gamma(\gamma-3)}{2}w_{\nu_{t}}^{2}-\gamma^{2}w_{\nu_{t}}^{2}$
$\geq$ $(\phi_{t}^{-1}\Delta_{\Sigma_{t}}\phi_{t}+\gamma u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\phi_{t}^{-1})+\frac{\gamma(\gamma-3)}{2}u^{-2}|\nabla_{\Sigma}u|^{2}$
$-\tfrac{1}{2}\operatorname{Sc}_{\Sigma}+\tilde{H}(t)\left(-w_{\nu_{t}}\right).$
Then
$-\phi_{t}^{-1}\left(\Delta_{\Sigma_{t}}\phi_{t}+\gamma\phi_{t}u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\right)-\frac{\gamma(\gamma-3)}{2}|\nabla_{\Sigma}w|^{2}$
$\geq$ $\phi_{t}^{-1}\tilde{H}^{\prime}(t)-\tilde{H}(t)w_{\nu_{t}}-\tfrac{1}{2}\operatorname{Sc}_{\Sigma}.$
Let $\phi_{t}=u^{-\frac{\gamma}{2}}e^{\xi_{t}}$ for some positive smooth function $\xi_{t}$ on $\Sigma_{t}$. Then
$\phi_{t}^{-1}\left(\Delta_{\Sigma_{t}}\phi_{t}+\gamma\phi_{t}u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\right)$
$=$ $\tfrac{\gamma(\gamma+2)}{4}u^{-2}|\nabla_{\Sigma}u|^{2}-\tfrac{\gamma}{2}u^{-1}\Delta_{\Sigma_{t}}u-\gamma\langle\nabla_{\Sigma}u,\nabla_{\Sigma_{t}}\xi_{t}\rangle u^{-1}$
$+\Delta_{\Sigma_{t}}\xi_{t}+|\nabla_{\Sigma_{t}}\xi_{t}|^{2}+\gamma u^{-1}\Delta_{\Sigma_{t}}u$
$-\tfrac{\gamma^{2}}{2}u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}u\rangle u^{-1}+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\xi_{t}\rangle$
$=$ $\tfrac{\gamma(2-\gamma)}{4}u^{-2}|\nabla_{\Sigma}u|^{2}+\tfrac{\gamma}{2}u^{-1}\Delta_{\Sigma_{t}}u+\Delta_{\Sigma_{t}}\xi_{t}+|\nabla_{\Sigma_{t}}\xi_{t}|^{2}$
$=$ $\tfrac{\gamma(4-\gamma)}{4}|\nabla_{\Sigma}w|^{2}+\tfrac{\gamma}{2}\Delta_{\Sigma_{t}}w+\Delta_{\Sigma_{t}}\xi_{t}+|\nabla_{\Sigma_{t}}\xi_{t}|^{2}.$
Hence
$-\phi_{t}^{-1}\left(\Delta_{\Sigma_{t}}\phi_{t}+\gamma\phi_{t}u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\right)-\frac{\gamma(\gamma-3)}{2}|\nabla_{\Sigma}w|^{2}$
$=$ $-\tfrac{\gamma(\gamma-2)}{4}|\nabla_{\Sigma}w|^{2}-\tfrac{\gamma}{2}\Delta_{\Sigma_{t}}w-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}.$
Therefore,
$\tilde{H}^{\prime}(t)-\tilde{H}(t)w_{\nu_{t}}\phi_{t}\leq\left(-\tfrac{\gamma(\gamma-2)}{4}|\nabla_{\Sigma}w|^{2}-\tfrac{\gamma}{2}\Delta_{\Sigma}w-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}+\tfrac{1}{2}\operatorname{Sc}_{\Sigma}\right)\phi_{t}.$
We can find a positive function $\varphi_{t}(x)\in C^{\infty}(\Sigma_{t})$ for $t\in(-\epsilon,\epsilon)$ such that
$\tilde{H}^{\prime}(t)-\tilde{H}(t)\frac{\int_{\Sigma_{t}}w_{\nu_{t}}\phi_{t}\varphi_{t}}{\int_{\Sigma_{t}}\varphi_{t}}$
$\leq$ $\frac{\int_{\Sigma_{t}}\left(-\tfrac{\gamma(\gamma-2)}{4}|\nabla_{\Sigma}w|^{2}-\tfrac{\gamma}{2}\Delta_{\Sigma}w-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}+\tfrac{1}{2}\operatorname{Sc}_{\Sigma}\right)\phi_{t}\varphi_{t}}{\int_{\Sigma_{t}}\varphi_{t}}$
$\leq$ $0.$
Otherwise, for any positive function $\varphi_{t}$ in $C^{\infty}(\Sigma_{t})$,
$\int_{\Sigma_{t}}\left(-\tfrac{\gamma(\gamma-2)}{4}|\nabla_{\Sigma}w|^{2}-\tfrac{\gamma}{2}\Delta_{\Sigma}w-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}+\tfrac{1}{2}\operatorname{Sc}_{\Sigma}\right)\phi_{t}\varphi_{t}>0.$
Let $\psi^{2}_{t}=\phi_{t}\varphi_{t}$, then
$\int_{\Sigma_{t}}\left(-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}\right)\psi^{2}_{t}$
$=$ $\int_{\Sigma_{t}}\left[-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}\psi^{2}_{t}-2\langle\nabla_{\Sigma_{t}}\psi,\psi\nabla_{\Sigma_{t}}\xi_{t}\rangle+\operatorname{div}_{\Sigma_{t}}\left(\psi^{2}\nabla_{\Sigma_{t}}\xi_{t}\right)\right]$
$\leq$ $\int_{\Sigma_{t}}|\nabla_{\Sigma_{t}}\psi_{t}|^{2}.$
We have
$\int_{\Sigma_{t}}\left(|\nabla_{\Sigma}\psi_{t}|^{2}+\tfrac{1}{2}\operatorname{Sc}_{\Sigma}\psi_{t}^{2}\right)+\int_{\Sigma}\gamma\psi_{t}\langle\nabla_{\Sigma_{t}}w,\nabla_{\Sigma_{t}}\psi_{t}\rangle+\int_{\Sigma_{t}}\tfrac{\gamma(2-\gamma)}{4}|\nabla_{\Sigma}w|^{2}>0.$
Then
(4.5) $\int_{\Sigma}\left[\frac{2(n-2)}{n-3}|\nabla_{\Sigma}\psi_{t}|^{2}+\frac{1}{2}\operatorname{Sc}_{\Sigma}\psi_{t}^{2}\right]>0,$
which is a contradiction with $\Sigma_{t}$ admitted no positive scalar curvature metric. Therefore, there exists $\varphi_{t}$ such that
$\tilde{H}^{\prime}(t)-\tilde{H}(t)\frac{\int_{\Sigma_{t}}w_{\nu_{t}}\phi_{t}\varphi_{t}}{\int_{\Sigma_{t}}\varphi_{t}}\leqslant 0.$
We set $\Psi(t)=-\frac{\int_{\Sigma_{t}}w_{\nu_{t}}\phi_{t}\varphi_{t}}{\int_{\Sigma_{t}}\varphi_{t}}$, then
$\tilde{H}^{\prime}+\Psi(t)\tilde{H}\leq 0.$
By solving this inequality, we finish the proof of the Proposition. ∎
Proposition 4.4.
Every $\Sigma_{t}$ is a stable warped area-minimizing hypersurface.
Proof.
Let $\partial_{t}$ be the variation vector field of the foliation $\{\Sigma_{t}\}_{t\in(-\varepsilon,\varepsilon)}$, and $\phi_{t}=\langle\partial_{t},\nu_{t}\rangle$. Recall that the first variation $E$, and
$\tfrac{\mathrm{d}}{\mathrm{d}t}E(\Sigma_{t})=\int_{\Sigma_{t}}\tilde{H}(t)u^{\gamma}\phi_{t}.$
It follows from $\tilde{H}(0)=0$ and (4.3) that $\tilde{H}(t)\leq 0$ for $t\geq 0$ and $\tilde{H}(t)\geq 0$ for $t\leq 0$. So $E(\Sigma_{t})\leq E(\Sigma_{0})$ for all $t\in(-\varepsilon,\varepsilon)$ and hence
$E(\Sigma_{t})=E(\Sigma_{0})$
for all $t\in(-\varepsilon,\varepsilon)$. Hence, all the foliation analysis on $\Sigma_{0}$ can be applied to $\Sigma_{t}$. Since $M$ is connected, we can conclude that $M$ is foliated by the stable warped area-minimizing hypersurface $\Sigma_{t}$. ∎
In the following, we deduce that the metric $g$ can be locally written as a warped product metric. From lemma 4.1 and lemma 4.2, we can obtain that, locally, $M^{n}$ can be written as $(-\epsilon,\epsilon)\times\Sigma$ such that each $\Sigma_{t}$ is a warped area-minimizing hypersurface and $\Sigma_{t}$ is umbilical. Then the metric $g$ can be written as $g=dt^{2}+f^{2}(t)g_{\Sigma}$ for some positive function $f$ on $(-\epsilon,\epsilon)$. Since $H_{t}=-(n-1)u^{-1}u^{\prime}(t)=(n-1)f^{-1}f^{\prime}(t)$, we obtain $f(t)=u^{-1}(t)$. By an open and closed argument, by passing to a suitable cover $\tilde{M}$, $(\tilde{M}^{n},g)$ is isometric to $(\mathbb{S}^{1}\times\Sigma^{n-1},ds^{2}+f^{2}(t)g|_{\Sigma})$. By the equation (1.3), we can obtain that $(\Sigma,g_{\Sigma})$ is scalar flat. Because $\Sigma$ admits no positive scalar curvature, we can deduce that $(\Sigma,g_{\Sigma})$ is Ricci flat.

### 21 Mar 2026 v2

@@
-warped area minimal hypersurface and warped product metric
+warped area-minimizing hypersurface and warped product metric
 Abstract.
@@
 1. Introduction
-A classic rigidity result concerning area-minimizing hypersurfaces, due to Cai-Galloway[CG00], asserts the following: if a three-dimensional manifold $(M,g)$ has non-negative scalar curvature ($\operatorname{Sc}_{g}\geq 0$) and contains a two-sided, locally area-minimizing torus $\Sigma$, then $M$ is flat in a neighborhood of $\Sigma$; equivalently, $M$ admits a product metric near $\Sigma$.
+For compact manifolds, a classic rigidity result concerning area-minimizing hypersurfaces, due to Cai-Galloway[CG00], asserts the following: if a three-dimensional manifold $(M,g)$ has non-negative scalar curvature ($\operatorname{Sc}_{g}\geq 0$) and contains a two-sided, locally area-minimizing torus $\Sigma$, then $M$ is flat in a neighborhood of $\Sigma$; equivalently, $M$ admits a product metric near $\Sigma$.
 This confirms a conjecture originally proposed by Fischer-Colbrie-Schoen[FCS80].
@@
 This implies that, in the rigid case, $M$ locally admits a product metric near $\Sigma$.
-Since then, additional rigidity results establishing local product metric splittings for area-minimizing hypersurfaces have been obtained [CEM19],[Zhu20],[CLZ25].
-A natural question thus arises: under what conditions do hypersurfaces admit local splittings as warped product metrics?
-In this paper, we address this question and establish several key results.
+Later, Zhu[Zhu20] extended Bray-Brendle-Neves’ result to a more general setting for an $n$-dimensional manifold $M^{n}$ that admits a non-zero degree map to $\mathbb{S}^{2}\times\mathbb{T}^{n}$, where $3\leq n+2\leq 7$.
+Recently, in the rigid case of their theorem, Chu-Lee-Zhu[CLZ25] proved that the universal cover of an $n$-dimensional manifold $(M^{n},g)$ with non-trivial homology group $H_{n}(M)$ splits isometrically as a product, assuming the bi-Ricci curvature is bounded from below by n1n-1.
+For noncompact manifolds, there is also an extensive literature concerning metric splitting theorems in the presence of area-minimizing hypersurfaces.
+For instance, we refer the reader to [AR89],[LW04],[CEM19],[CCE16],[HSY25],[Zhu23].
+A natural question thus arises: under what conditions do hypersurfaces admit local splitting as warped product metrics?
+In this paper, we address this question for compact manifolds and establish several key results.
 We further elucidate the geometric significance of warped(or weighted) area-minimizing hypersurfaces.
@@
 is constant. Hence, the first and second variations of $E(N^{n-1}_{t})$ along the $\partial_{t}$ direction are all zero.
+The equations (1.4) and (1.5) are similar to the spectral curvature.
 Therefore, we study the warped minimal hypersurface under the spectral Ricci curvature condition and the spectral scalar curvature condition.
@@
 is the least Ricci curvature at a given point $x\in M$.
-Throughout this paper, we let $\gamma=n-1$.
+Spectral curvature plays an important role in the study of the stable Bernstein problem; see [CLMS24],[Maz24].
+A variety of geometric results have been established concerning spectral curvature, including width estimates [HKKZ24],[CS25], diameter and volume estimates [CR21], [AX24],[CH26],[Wu26] and splitting theorems(product case) [APX24],[CMMR25],[HW26],[Yeu25].
+We obtain some splitting theorems(warped product case) under the spectral curvature condition as follows. Throughout this paper, we let $\gamma=n-1$.
 Theorem 1.2.
@@
 As Antonelli-Pozzetta-Xu pointed out, such a product metric splitting fails to hold when $a\geq\frac{4}{n-1}$.
-Nevertheless, for the case $n=3$, if there exists a compact warped area-minimizing hypersurface, then the manifold $M^{3}$ with two ends will locally split as a warped product metric according to Theorem 1.2.
+Nevertheless, for the case $n=3$ and a=2, if there exists a compact warped area-minimizing hypersurface, then the manifold $M^{3}$ with two ends will locally split as a warped product metric according to Theorem 1.2.
 For nonnegative Bakry–Émery Ricci curvature, Liu [Liu13] obtains a locally split product metric by studying the weighted area-minimizing hypersurface.
+Li-Wang[LW06] also proved that the manifold splits as a warped product under a lower Ricci curvature bound and additional geometric conditions.
 For the spectral scalar curvature, we have
@@
 (1.10) $\gamma u^{-1}\Delta_{g}u+\frac{1}{2}\operatorname{Sc}_{g}\geq\frac{\gamma(\gamma-3)}{2}u^{-2}|\nabla u|^{2},$
-where $u$ is a smooth positive function on $M$ and $5\leq n\leq 7$.
+where $u$ is a smooth positive function on $M$ and $4\leq n\leq 7$.
 If $H_{n-1}(M,\mathbb{Z})\neq 0$ and $0\neq[\Sigma]\in H_{n-1}(M,\mathbb{Z})$ such that there is no positive scalar curvature on $\Sigma$, then, $u$ is constant on $\Sigma$ and $(M^{n},g)$ locally splits as $(M^{n},g=dt^{2}+u^{-2}g_{\Sigma})$.
@@
 Remark 1.5.
-The operator $\gamma\Delta_{g}+\frac{1}{2}\operatorname{Sc}_{g}$ is greater than the conformal operator $\frac{2(n-1)}{n-2}\Delta_{g}+\frac{1}{2}\operatorname{Sc}_{g}$ for $5\leq n\leq 7$.
-That is, if the $\gamma\Delta_{g}+\frac{1}{2}\operatorname{Sc}_{g}$ is positive, we can not use the conformal change to obtain a positive scalar curvature metric.
+If we let $v=u^{\frac{n-2}{2}}$, then the inequality (1.10) implies that
+$\frac{2(n-1)}{n-2}v^{-1}\Delta_{g}v+\frac{1}{2}\operatorname{Sc}_{g}\geq 0.$
+Then we can use the conformal change to obtain a nonnegative scalar curvature metric.
 The proofs of Theorem 1.2 and Theorem 1.4 rely on the study of warped(or weighted) area-minimizing hypersurfaces.
@@
 In section 4, we prove Theorem 1.4.
+Acknowledgements: The author would like to express his gratitude to Professor Thomas Richard and Professor Hong Han for their interest and conversation about this work.
 2. Basic fact about warped or weighted minimal hypersurface
@@
 where $\gamma=n-1$ and $\mathcal{H}^{n-1}$ denotes the (n1)(n-1) dimension Hausdorff measure induced by the metric $g$.
-Let $H$ and $A$ stand for the mean curvature and the second fundamental of $\Sigma$, respectively.
-Lemma 2.1.
+Let $H$ and AA stand for the mean curvature and the second fundamental of $\Sigma$, respectively.
+First, we give the definition of the warped area-minimizing hypersurface.
+Definition 2.1.
+Let $\Sigma$ be a compact hypersurface in a Riemannian manifold $(M,g)$.
+Suppose that the homology class $\Sigma]\in H_{n-1}(M,\mathbb{Z})$ of $\Sigma$ is non-zero, i.e., $\Sigma]\neq 0$ in $H_{n-1}(M,\mathbb{Z})$.
+Then $\Sigma$ is called a warped area-minimizing hypersurface if
+$\int_{\Sigma}u^{\gamma}\mathrm{d}\mathcal{H}^{n-1}=\inf_{\hat{\Sigma}\in[\Sigma]}\{\int_{\hat{\Sigma}}u^{\gamma}\mathrm{d}\mathcal{H}^{n-1}\}.$
+Lemma $2.$
 Let $\Sigma_{t}$ be a variation of $\Sigma$ with the variation vector field given by $Y$ and $\nu$ be the unit normal vector of $\Sigma$, then
 (2.2) $\tfrac{\mathrm{d}}{\mathrm{d}t}E(\Sigma_{t})|_{t=0}=\int_{\Sigma}(H+\gamma u^{-1}u_{\nu})\langle Y,\nu\rangle u^{\gamma}\mathrm{d}\mathcal{H}^{n-1}.$
-We say $\Sigma$ a warped(or weighted) area-minimizing hypersurface if $H+\gamma u^{-1}u_{\nu}=0$ along $\Sigma$.
+We say $\Sigma$ a warped(or weighted) minimal hypersurface if $H+\gamma u^{-1}u_{\nu}=0$ along $\Sigma$.
 Lemma 2.3.
-Consider a warped area-minimizing hypersurface $\Sigma$.
-Assume that $\Sigma_{t}$ is a smooth $1$-parameter family of smooth hypersurfaces along the variation vector field $Y$ and normal speed $\phi=\langle Y,\nu\rangle$ at $t=0$.
+Consider a warped minimal hypersurface $\Sigma$.
+Assume that $\Sigma_{t}$ is a smooth 11-parameter family of smooth hypersurfaces along the variation vector field $Y$ and normal speed $\phi=\langle Y,\nu\rangle$ at $t=0$.
 Then
@@
 (2.4) $\qquad+\gamma u^{-1}\phi(\Delta u-\Delta_{\partial\Omega}u-Hu_{\nu})-\gamma u^{-1}\langle\nabla_{\partial\Omega}u,\nabla_{\partial\Omega}\phi\rangle]u^{\gamma}\phi.$
-A warped area-minimizing hypersurface $\Sigma$ is called stable if
+A warped minimal hypersurface $\Sigma$ is called stable if
 $\tfrac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}E(\Sigma_{t})|_{t=0}\geq 0.$
@@
 Then the stable warped area-minimizing property of $\Sigma_{t}$ follows.
-We do that in Proposition 4.3 and $4.$ ∎
+We do that in Proposition 4.3 and 4.4.
+∎
 Proposition 4.3.
@@
 $\quad-\gamma u^{-1}(\Delta_{g}u-\Delta_{\Sigma_{t}}u-H_{t}u_{\nu_{t}})\phi_{t}+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle$
-Using the spectral scalar curvature bound, and with suitable grouping of terms, we see
+Using the spectral scalar curvature bound and (4.4) with suitable grouping of terms, we see
 $-\tilde{H}^{\prime}(t)\phi_{t}^{-1}$
-$\geq$ $(\phi_{t}^{-1}\Delta_{\Sigma_{t}}\phi_{t}+\gamma u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\phi_{t}^{-1})-\tfrac{1}{2}\operatorname{Sc}_{\Sigma}$
+$\geq$ $(\phi_{t}^{-1}\Delta_{\Sigma_{t}}\phi_{t}+\gamma u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\phi_{t}^{-1})-\tfrac{1}{2}\operatorname{Sc}_{\Sigma_{t}}$
 $\quad+\frac{n}{2(n-1)}H_{t}^{2}+\gamma w_{\nu_{t}}^{2}+\gamma H_{t}w_{\nu_{t}}+\frac{\gamma(\gamma-3)}{2}u^{-2}|\nabla u|^{2}$
-$\geq$ $(\phi_{t}^{-1}\Delta_{\Sigma_{t}}\phi_{t}+\gamma u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\phi_{t}^{-1})+\frac{\gamma(\gamma-3)}{2}u^{-2}|\nabla_{\Sigma}u|^{2}$
-$\quad-\tfrac{1}{2}\operatorname{Sc}_{\Sigma}+\frac{n}{2(n-1)}\tilde{H}^{2}(t)+\tilde{H}(t)\left(\gamma w_{\nu_{t}}-\frac{n\gamma}{n-1}w_{\nu_{t}}\right)$
+$\geq$ $(\phi_{t}^{-1}\Delta_{\Sigma_{t}}\phi_{t}+\gamma u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\phi_{t}^{-1})+\frac{\gamma(\gamma-3)}{2}u^{-2}|\nabla_{\Sigma_{t}}u|^{2}$
+$\quad-\tfrac{1}{2}\operatorname{Sc}_{\Sigma_{t}}+\frac{n}{2(n-1)}\tilde{H}^{2}(t)+\tilde{H}(t)\left(\gamma w_{\nu_{t}}-\frac{n\gamma}{n-1}w_{\nu_{t}}\right)$
 $+\gamma w_{\nu_{t}}^{2}+\frac{n\gamma^{2}}{2(n-1)}w_{\nu_{t}}^{2}+\frac{\gamma(\gamma-3)}{2}w_{\nu_{t}}^{2}-\gamma^{2}w_{\nu_{t}}^{2}$
-$\geq$ $(\phi_{t}^{-1}\Delta_{\Sigma_{t}}\phi_{t}+\gamma u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\phi_{t}^{-1})+\frac{\gamma(\gamma-3)}{2}u^{-2}|\nabla_{\Sigma}u|^{2}$
-$-\tfrac{1}{2}\operatorname{Sc}_{\Sigma}+\tilde{H}(t)\left(-w_{\nu_{t}}\right).$
+$\geq$ $(\phi_{t}^{-1}\Delta_{\Sigma_{t}}\phi_{t}+\gamma u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\phi_{t}^{-1})+\frac{\gamma(\gamma-3)}{2}u^{-2}|\nabla_{\Sigma_{t}}u|^{2}$
+$-\tfrac{1}{2}\operatorname{Sc}_{\Sigma_{t}}+\tilde{H}(t)\left(-w_{\nu_{t}}\right).$
 Then
 $-\phi_{t}^{-1}\left(\Delta_{\Sigma_{t}}\phi_{t}+\gamma\phi_{t}u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\right)-\frac{\gamma(\gamma-3)}{2}|\nabla_{\Sigma}w|^{2}$
-$\geq$ $\phi_{t}^{-1}\tilde{H}^{\prime}(t)-\tilde{H}(t)w_{\nu_{t}}-\tfrac{1}{2}\operatorname{Sc}_{\Sigma}.$
+$\geq$ $\phi_{t}^{-1}\tilde{H}^{\prime}(t)-\tilde{H}(t)w_{\nu_{t}}-\tfrac{1}{2}\operatorname{Sc}_{\Sigma_{t}}.$
 Let $\phi_{t}=u^{-\frac{\gamma}{2}}e^{\xi_{t}}$ for some positive smooth function $\xi_{t}$ on $\Sigma_{t}$. Then
 $\phi_{t}^{-1}\left(\Delta_{\Sigma_{t}}\phi_{t}+\gamma\phi_{t}u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\right)$
-$=$ $\tfrac{\gamma(\gamma+2)}{4}u^{-2}|\nabla_{\Sigma}u|^{2}-\tfrac{\gamma}{2}u^{-1}\Delta_{\Sigma_{t}}u-\gamma\langle\nabla_{\Sigma}u,\nabla_{\Sigma_{t}}\xi_{t}\rangle u^{-1}$
+$=$ $\tfrac{\gamma(\gamma+2)}{4}u^{-2}|\nabla_{\Sigma_{t}}u|^{2}-\tfrac{\gamma}{2}u^{-1}\Delta_{\Sigma_{t}}u-\gamma\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\xi_{t}\rangle u^{-1}$
 $+\Delta_{\Sigma_{t}}\xi_{t}+|\nabla_{\Sigma_{t}}\xi_{t}|^{2}+\gamma u^{-1}\Delta_{\Sigma_{t}}u$
 $-\tfrac{\gamma^{2}}{2}u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}u\rangle u^{-1}+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\xi_{t}\rangle$
-$=$ $\tfrac{\gamma(2-\gamma)}{4}u^{-2}|\nabla_{\Sigma}u|^{2}+\tfrac{\gamma}{2}u^{-1}\Delta_{\Sigma_{t}}u+\Delta_{\Sigma_{t}}\xi_{t}+|\nabla_{\Sigma_{t}}\xi_{t}|^{2}$
-$=$ $\tfrac{\gamma(4-\gamma)}{4}|\nabla_{\Sigma}w|^{2}+\tfrac{\gamma}{2}\Delta_{\Sigma_{t}}w+\Delta_{\Sigma_{t}}\xi_{t}+|\nabla_{\Sigma_{t}}\xi_{t}|^{2}.$
+$=$ $\tfrac{\gamma(2-\gamma)}{4}u^{-2}|\nabla_{\Sigma_{t}}u|^{2}+\tfrac{\gamma}{2}u^{-1}\Delta_{\Sigma_{t}}u+\Delta_{\Sigma_{t}}\xi_{t}+|\nabla_{\Sigma_{t}}\xi_{t}|^{2}$
+$=$ $\tfrac{\gamma(4-\gamma)}{4}|\nabla_{\Sigma_{t}}w|^{2}+\tfrac{\gamma}{2}\Delta_{\Sigma_{t}}w+\Delta_{\Sigma_{t}}\xi_{t}+|\nabla_{\Sigma_{t}}\xi_{t}|^{2}.$
 Hence
-$-\phi_{t}^{-1}\left(\Delta_{\Sigma_{t}}\phi_{t}+\gamma\phi_{t}u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\right)-\frac{\gamma(\gamma-3)}{2}|\nabla_{\Sigma}w|^{2}$
-$=$ $-\tfrac{\gamma(\gamma-2)}{4}|\nabla_{\Sigma}w|^{2}-\tfrac{\gamma}{2}\Delta_{\Sigma_{t}}w-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}.$
+$-\phi_{t}^{-1}\left(\Delta_{\Sigma_{t}}\phi_{t}+\gamma\phi_{t}u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\right)-\frac{\gamma(\gamma-3)}{2}|\nabla_{\Sigma_{t}}w|^{2}$
+$=$ $-\tfrac{\gamma(\gamma-2)}{4}|\nabla_{\Sigma_{t}}w|^{2}-\tfrac{\gamma}{2}\Delta_{\Sigma_{t}}w-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}.$
 Therefore,
-$\tilde{H}^{\prime}(t)-\tilde{H}(t)w_{\nu_{t}}\phi_{t}\leq\left(-\tfrac{\gamma(\gamma-2)}{4}|\nabla_{\Sigma}w|^{2}-\tfrac{\gamma}{2}\Delta_{\Sigma}w-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}+\tfrac{1}{2}\operatorname{Sc}_{\Sigma}\right)\phi_{t}.$
+$\tilde{H}^{\prime}(t)-\tilde{H}(t)w_{\nu_{t}}\phi_{t}\leq\left(-\tfrac{\gamma(\gamma-2)}{4}|\nabla_{\Sigma_{t}}w|^{2}-\tfrac{\gamma}{2}\Delta_{\Sigma_{t}}w-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}+\tfrac{1}{2}\operatorname{Sc}_{\Sigma_{t}}\right)\phi_{t}.$
 We can find a positive function $\varphi_{t}(x)\in C^{\infty}(\Sigma_{t})$ for $t\in(-\epsilon,\epsilon)$ such that
 $\tilde{H}^{\prime}(t)-\tilde{H}(t)\frac{\int_{\Sigma_{t}}w_{\nu_{t}}\phi_{t}\varphi_{t}}{\int_{\Sigma_{t}}\varphi_{t}}$
-$\leq$ $\frac{\int_{\Sigma_{t}}\left(-\tfrac{\gamma(\gamma-2)}{4}|\nabla_{\Sigma}w|^{2}-\tfrac{\gamma}{2}\Delta_{\Sigma}w-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}+\tfrac{1}{2}\operatorname{Sc}_{\Sigma}\right)\phi_{t}\varphi_{t}}{\int_{\Sigma_{t}}\varphi_{t}}$
+$\leq$ $\frac{\int_{\Sigma_{t}}\left(-\tfrac{\gamma(\gamma-2)}{4}|\nabla_{\Sigma_{t}}w|^{2}-\tfrac{\gamma}{2}\Delta_{\Sigma_{t}}w-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}+\tfrac{1}{2}\operatorname{Sc}_{\Sigma_{t}}\right)\phi_{t}\varphi_{t}}{\int_{\Sigma_{t}}\varphi_{t}}$
 $\leq$ $0.$
 Otherwise, for any positive function $\varphi_{t}$ in $C^{\infty}(\Sigma_{t})$,
-$\int_{\Sigma_{t}}\left(-\tfrac{\gamma(\gamma-2)}{4}|\nabla_{\Sigma}w|^{2}-\tfrac{\gamma}{2}\Delta_{\Sigma}w-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}+\tfrac{1}{2}\operatorname{Sc}_{\Sigma}\right)\phi_{t}\varphi_{t}>0.$
+$\int_{\Sigma_{t}}\left(-\tfrac{\gamma(\gamma-2)}{4}|\nabla_{\Sigma_{t}}w|^{2}-\tfrac{\gamma}{2}\Delta_{\Sigma_{t}}w-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}+\tfrac{1}{2}\operatorname{Sc}_{\Sigma_{t}}\right)\phi_{t}\varphi_{t}>0.$
 Let $\psi^{2}_{t}=\phi_{t}\varphi_{t}$, then
@@
 We have
-$\int_{\Sigma_{t}}\left(|\nabla_{\Sigma}\psi_{t}|^{2}+\tfrac{1}{2}\operatorname{Sc}_{\Sigma}\psi_{t}^{2}\right)+\int_{\Sigma}\gamma\psi_{t}\langle\nabla_{\Sigma_{t}}w,\nabla_{\Sigma_{t}}\psi_{t}\rangle+\int_{\Sigma_{t}}\tfrac{\gamma(2-\gamma)}{4}|\nabla_{\Sigma}w|^{2}>0.$
+$\int_{\Sigma_{t}}\left(|\nabla_{\Sigma_{t}}\psi_{t}|^{2}+\tfrac{1}{2}\operatorname{Sc}_{\Sigma}\psi_{t}^{2}\right)+\int_{\Sigma}\gamma\psi_{t}\langle\nabla_{\Sigma_{t}}w,\nabla_{\Sigma_{t}}\psi_{t}\rangle+\int_{\Sigma_{t}}\tfrac{\gamma(2-\gamma)}{4}|\nabla_{\Sigma_{t}}w|^{2}>0.$
 Then
-(4.5) $\int_{\Sigma}\left[\frac{2(n-2)}{n-3}|\nabla_{\Sigma}\psi_{t}|^{2}+\frac{1}{2}\operatorname{Sc}_{\Sigma}\psi_{t}^{2}\right]>0,$
+(4.5) $\int_{\Sigma}\left[\frac{2(n-2)}{n-3}|\nabla_{\Sigma_{t}}\psi_{t}|^{2}+\frac{1}{2}\operatorname{Sc}_{\Sigma_{t}}\psi_{t}^{2}\right]>0,$
 which is a contradiction with $\Sigma_{t}$ admitted no positive scalar curvature metric. Therefore, there exists $\varphi_{t}$ such that
@@
 By an open and closed argument, by passing to a suitable cover $\tilde{M}$, $(\tilde{M}^{n},g)$ is isometric to $(\mathbb{S}^{1}\times\Sigma^{n-1},ds^{2}+f^{2}(t)g|_{\Sigma})$.
-By the equation (1.3), we can obtain that $(\Sigma,g_{\Sigma})$ is scalar flat.
-Because $\Sigma$ admits no positive scalar curvature, we can deduce that $(\Sigma,g_{\Sigma})$ is Ricci flat.
+Because $(\Sigma,g_{\Sigma})$ is scalar flat and $\Sigma$ admits no positive scalar curvature, we can deduce that $(\Sigma,g_{\Sigma})$ is Ricci flat.